OpenStudy (anonymous):

how to find the range and domain of the following: f(x)=sqr root of (1-3x)

OpenStudy (anonymous):

domain: solve \(1-3x\geq 0\) for \(x\)

OpenStudy (anonymous):

range is \([0\infty)\) or \(f\geq 0\) since the radical means the positive square root

OpenStudy (anonymous):

can you solve \(1-3x\geq0\) ?

OpenStudy (anonymous):

yup...thnx... :) so, to find range of sqr root, there must be no -ve ?

OpenStudy (anonymous):

yes \[\sqrt{whatever}\geq 0\] always for example \(\sqrt{9}=3,\sqrt0=0\) it is never negative

OpenStudy (anonymous):

ok... what about to find range of

OpenStudy (anonymous):

\[\sqrt{16-x ^{2}}\]